# Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger

### v1

@@ C.1 Prompting Strategies
    CoT (Chain of Thought): The model is instructed to think step-by-step, reasoning why it does or does not need external tools to address the user query, and finally concludes its decision with “Yes" or “No."
[…]
Note that there are no results for the CoT prompting strategy for the Mistral-7b-instruct-v0.3 model. Regardless of the prompts used, the model consistently responds with “Yes/No" at the beginning, followed by an explanation of its decision. This behavior effectively mirrors the Yes/No+Explanation prompting strategy. Based on Table 6, we make the following observations and provide corresponding analysis:
@@
    Yes/No prompting strategy works better than No/Yes prompting. We hypothesize this phenomenon is due to the data format in the pre-training data. For example, there are likely many more Yes/No answers and reasoning processes in the training data compared to No/Yes answers, influencing the model’s performance.
@@
Table 6: Performance Comparison of Different Prompting Strategies
@@ C.2 P(Yes) Approach
The $P_{\text{Yes}}$ baseline introduces a Yes-score, as defined in Section 5. This score provides a nuanced measure of the model’s confidence, refining the binary approach taken by the Naive baseline. The Yes-score spans from 00 to $1$, where a score of 00 signifies a definite "No" and a score of $1$ signifies a definite “Yes". Scores close to $0.5$ reflect lower certainty in the model’s response, signifying ambiguity in decision-making. By adjusting the model’s output in cases where the Yes-score is near $0.5$ to always “Yes/No" answer, we aim to enhance the accuracy of both tool use and RAG timing. We employ 2 to determine the optimal threshold $l$ for the Yes-score based on training data, which is then applied to the test data.
@@ C.3 Distribution of P(Yes) and Meta-Cognition Scores
In Figure LABEL:fig:p_yes_distribution, we present the distribution of P(Yes) scores for both correct and incorrect Yes/No decisions. Our key observations are as follows:
@@
    When the model is given detailed instructions and few-shot examples, it demonstrates poor calibration. As illustrated in Figure LABEL:fig:p_yes_distribution(a), the distributions of P(Yes) scores for correct and incorrect decisions do not show a clear distinction.
@@
    After fine-tuning, the model displays significantly better calibration, as shown in Figures LABEL:fig:p_yes_distribution(c) and (d). Most correct decisions have P(Yes) scores of either 1 (indicating "Yes") or 0 (indicating "No"), while the P(Yes) scores for incorrect decisions vary between 0 and 1.
@@ D.2 Different Size of Training Data
We further examine how the size of the training data affects the outcomes of the meta-cognition probe. Specifically, we analyze the performance of the trained probes with varying sizes of training data, as illustrated in Figure 16 and Figure LABEL:fig:probe_n_train_rag. According to Equation 1, a sentence with 10 tokens can be used to create 10 training data pairs of experimental prompts and reference prompts. Typically, a brief explanation of why or why not to use external tools/RAG corresponds to around 30 to 50 tokens. Thus, a training data size of 256 requires fewer than 10 queries and their associated explanations.
@@ D.3 Contrastive Instructions for Various Probes
[…]
@@
[…]
@@
[…]
@@
[…]
@@ E.1 Prompts in Adaptive Tool Use
Prompt with context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction If you think it’s necessary to use external tools, please respond with “Yes"; otherwise, respond with “No". Additionally, you should provide a very brief explanation for your answer. Here are some examples: Query: “Write an opinion piece about why diversity and inclusion is super important for the tech industry. The essay should be targeted at ’tech bros’, and should avoid alienating them, but instead appeal to their logic; it should explain how diversity and inclusion of women, immigrants, etc. could benefit them specifically." Answer: No Query: “Are there any loopholes that hackers can exploit on my website?" Answer: Yes Query: “Plan a weekly lunch menu for a school. Write down a main dish, a carbohydrate side dish, a vegetable side dish, and a dessert for each day." Answer: No Query: “Can you break down the main points of this TED talk for me? Here’s the YouTube link." Answer: Yes Query: “How’s the weather in London right now?" Answer: No
[…]
@@ E.2 Prompts in Adaptive RAG
[…]

### v2

@@ C.1 Prompting Strategies
    CoT (Chain of Thought): The model is instructed to think step-by-step, reasoning why it does or does not need external tools to address the user query, and finally concludes its decision with "Yes" or "No."
Note that there are no results for the CoT prompting strategy for the Mistral-7b-instruct-v0.3 model. Regardless of the prompts used, the model consistently responds with "Yes/No" at the beginning, followed by an explanation of its decision. This behavior effectively mirrors the Yes/No+Explanation prompting strategy. Based on Table 8, we make the following observations and provide corresponding analysis:
@@
    Yes/No prompting strategy works better than No/Yes prompting. We hypothesize that this phenomenon is due to the data format in the pre-training data. For example, there are likely many more Yes/No answers and reasoning processes in the training data compared to No/Yes answers, influencing the model’s performance.
Chain of Thought Prompting. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction Please think step by step, and provide a brief explanation for your decision at first. At last, please conclude with "Yes" if you need to use external tools, or "No" if you do not need external tools. {Few-shot Examples}
User query: {query}
Answer:
@@
Table 8: Performance comparison of different prompting strategies.
@@ C.2 P(Yes) Approach
The $P_{\text{Yes}}$ baseline introduces a Yes-score, as defined in Section 5. This score provides a nuanced measure of the model’s confidence, refining the binary approach taken by the Naive baseline. The Yes-score spans from 0 to $1$, where a score of 0 signifies a definite "No" and a score of $1$ signifies a definite "Yes". Scores close to $0.5$ reflect lower certainty in the model’s response, signifying ambiguity in decision-making. By adjusting the model’s output in cases where the Yes-score is near $0.5$ to always "Yes/No" answer, we aim to enhance the accuracy of both tool use and RAG timing. We employ Equation (2) to determine the optimal threshold $l$ for the Yes-score based on training data, which is then applied to the test data.
@@ C.3 Distribution of P(Yes) and Meta-Cognition Scores
In Figure LABEL:fig:p_yes_distribution, we present the distribution of $P_{\text{Yes}}$ scores for both correct and incorrect Yes/No decisions. Our key observations are as follows:
@@
    When the model is given detailed instructions and few-shot examples, it demonstrates poor calibration. As illustrated in Figure LABEL:fig:p_yes_distribution(a), the distributions of $P_{\text{Yes}}$ scores for correct and incorrect decisions do not show a clear distinction.
@@
    After fine-tuning, the model displays significantly better calibration, as shown in Figures LABEL:fig:p_yes_distribution(c) and (d). Most correct decisions have $P_{\text{Yes}}$ scores of either 1 (indicating "Yes") or 0 (indicating "No"), while the $P_{\text{Yes}}$ scores for incorrect decisions vary between 0 and 1.
@@ D.2 Different Size of Training Data
We further examine how the size of the training data affects the outcomes of the meta-cognition probe. Specifically, we analyze the performance of the trained probes with varying sizes of training data, as illustrated in Figure 15 and Figure LABEL:fig:probe_n_train_rag. According to Equation (1), a sentence with 10 tokens can be used to create 10 training data pairs of experimental prompts and reference prompts. Typically, a brief explanation of why or why not to use external tools/RAG corresponds to around 30 to 50 tokens. Thus, a training data size of 256 requires fewer than 10 queries and their associated explanations.
@@ D.3 Contrastive Instructions for Various Probes
Contrastive Instructions for training Honesty Probe Honest Instruction Imagine you are an honest intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Untruthful Instruction Imagine you are an untruthful intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
@@
Contrastive Instructions for training Confidence Probe Confident Instruction Imagine you are a confident intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Unconfident Instruction Imagine you are an unconfident intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
@@
Contrastive Instructions for training Meta-Cognition Probe Strong Meta-Cognition Instruction in Adaptive Tool Use Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. You can accurately assess and explain whether you need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Weak Meta-Cognition Instruction Imagine you are an assistant with weak meta-cognition, often unaware of your own limitations and capabilities. You struggle to assess and explain why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
@@
Contrastive Instructions for training Meta-Cognition Probe in Adaptive RAG Strong Meta-Cognition Instruction Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. You can accurately assess and explain whether you need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. User Query: {query}
Answer:
Weak Meta-Cognition Instruction Imagine you are an assistant with weak meta-cognition, often unaware of your own limitations and capabilities. You struggle to assess and explain why you need or do not need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. User Query: {query}
Answer:
@@ E.1 Prompts in Adaptive Tool Use
Prompt with context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction If you think it’s necessary to use external tools, please respond with "Yes"; otherwise, respond with "No". Additionally, you should provide a very brief explanation for your answer. Here are some examples: Query: "Write an opinion piece about why diversity and inclusion is super important for the tech industry. The essay should be targeted at ’tech bros’, and should avoid alienating them, but instead appeal to their logic; it should explain how diversity and inclusion of women, immigrants, etc. could benefit them specifically." Answer: No Query: "Are there any loopholes that hackers can exploit on my website?" Answer: Yes Query: "Plan a weekly lunch menu for a school. Write down a main dish, a carbohydrate side dish, a vegetable side dish, and a dessert for each day." Answer: No Query: "Can you break down the main points of this TED talk for me? Here’s the YouTube link." Answer: Yes Query: "How’s the weather in London right now?" Answer: No
User query: {query}
Answer:
Prompt without context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. If you think it’s necessary to use external tools, please respond with "Yes"; otherwise, respond with "No". Additionally, you should provide a very brief explanation for your answer. User Query: {query}
Answer:
@@ E.2 Prompts in Adaptive RAG
Prompt without context. Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. Your task is to accurately assess and explain whether you need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. If you determine that performing RAG is necessary, please respond with "Yes"; otherwise, respond with "No". Additionally, provide a very brief explanation for your decision. User Query: {query}
Answer: